OpenStudy (yanasidlinskiy):

Can someone help me with this question??Please..I don't know what to do next.

OpenStudy (yanasidlinskiy):

Multiply using Napier's rods. 3 x 4565 3*4 = 12 3*5 = 15 3*6 = 18 3*5 = 15

OpenStudy (taylorterminate):

@YanaSidlinskiy so, whats 3*4?

OpenStudy (yanasidlinskiy):

I already have it ^. Look.

OpenStudy (taylorterminate):

XD lol

OpenStudy (taylorterminate):

111 258 and, 3*5 is 15 again sooo 1111 2585 add em up

OpenStudy (yanasidlinskiy):

Add the 1111 and 2585 together?

OpenStudy (taylorterminate):

yes

OpenStudy (nurali):

just multiply each number by the 3 and stagger the results 1111 2585 ------- 13695 this is the principle behind the rods to begin with

OpenStudy (taylorterminate):

>_>

OpenStudy (yanasidlinskiy):

I got 3696. That doesn't seem right. I don't know what I'm doing wrong.

OpenStudy (taylorterminate):

try 2 use a calculator >_>

OpenStudy (yanasidlinskiy):

I used my calculator.

OpenStudy (taylorterminate):

do u have answer choices?

OpenStudy (yanasidlinskiy):

Nope!

OpenStudy (taylorterminate):

/).(\ @Nurali help please

OpenStudy (yanasidlinskiy):

Nurali, am I supposed to multiply by 3? And then add them? I don't understand.

OpenStudy (taylorterminate):

Ooooo i think ur answer would b http://prntscr.com/48gtm4

OpenStudy (yanasidlinskiy):

Ok. But I want to know how to get the answer.

OpenStudy (taylorterminate):

yes @Nurali just explained it 2 u

OpenStudy (yanasidlinskiy):

OK. I think I got it. Not sure yet though. But thanx anyways:D

OpenStudy (nurali):

https://www.youtube.com/watch?v=WZLpqTyZMM4

OpenStudy (taylorterminate):

now u come 2 help @Nurali /).(\